### v1

@@ Abstract
Electric vehicle charging demand prediction is important for vacant charging pile recommendation and charging infrastructure planning, thus facilitating vehicle electrification and green energy development. The performance of previous spatio-temporal studies is still far from satisfactory because the traditional graphs are difficult to model non-pairwise spatial relationships and multivariate temporal features are not adequately taken into account. To tackle these issues, we propose an attention-based heterogeneous multivariate data fusion approach (AHMDF) for citywide electric vehicle charging demand prediction, which incorporates geo-based clustered hypergraph and multivariate gated Transformer to considers both static and dynamic influences. To learn non-pairwise relationships, we cluster service areas by the types and numbers of points of interest in the areas and develop attentive hypergraph networks accordingly. Graph attention mechanisms are used for information propagation between neighboring areas. Additionally, we improve the Transformer encoder utilizing gated mechanisms so that it can selectively learn dynamic auxiliary information and temporal features. Experiments on an electric vehicle charging benchmark dataset demonstrate the effectiveness of our proposed approach compared with a broad range of competing baselines. Furthermore, we demonstrate the impact of dynamic influences on prediction results in different areas of the city and the effectiveness of our clustering method.
@@ 1 Introduction
As the world accelerates its transition to low-carbon and environmental friendly mobility, electric vehicles (EVs) will play an important role in the future automotive market [1]. The International Energy Agency expects that policies in major economies around the world will increase the market share of electric vehicles to 50% by 2030. However, the lack of available and reasonably located charging facilities is a major constraint to the development of electric vehicles [2, 3]. The increasing development of smartphones, in-vehicle devices and information distribution systems makes governments and mangers glad to turn to accurate EV charging demand prediction for help. By predicting the demand for services in the coming future, administrators can provide users with suggestive information, such as timing and route planning, to avoid customer losses. For governments, accurate charging demand prediction, especially at the citywide level, is a groundwork for a variety of smart strategies such as charging infrastructure planning and smart power supply [4, 5, 6].
Researchers have made many efforts to enable highly accurate EV charging demand prediction. First, due to the simplicity and computational speed of the statistical method, it has been widely used in early studies to regress time-series charging demand data, but the linear assumption limits its accuracy [7, 8]. With the development of computer technology, machine learning has become an effective method for charging demand prediction, especially spatio-temporal prediction, as it can accurately learn data patterns and fuse multidimensional features. Many models introduce Recurrent neural networks to extract temporal features [9, 10] and Graph neural networks to extract spatial features [11], respectively. However, the ambiguity in the weight assignment of convolutional and recurrent operations limits their capabilities. In the latest researches, time-series attention networks (e.g., Transformer) [12] and graph attention networks [13] that are able to better determine the weights on the network have become popular tools for improving the spatio-temporal prediction accuracy. However, classical attention mechanisms focus on the entire sequence of input hidden states, whereas not all hidden states can provide enough useful information for prediction, especially in long time series. Moreover, some typical models lack consideration of external influences and are difficult to cope with complex dynamics, making prediction lagging and less accurate [14].
@@
EV charging demand is typically influenced by many factors that can be categorized into two main groups, namely static and dynamic, as shown in Fig. 1. Static influences include neighborhood charging services and land use around the charging stations. These influences are heterogeneous to the time-series charging demands and are usually not directly acceptable as inputs to the network. For example, improvements of the road network and the smart application on mobile devices make it possible for demand in neighboring areas to influence each other. Previous studies employ graph neural networks to fuse information from adjacent nodes and achieve good results [15, 16]. Moreover, some areas have similar land uses and residents may have typical and similar charging habits within those areas [17]. For example, charging demand in commercial areas is relatively high on weekends and after work on weekdays, while it’s the opposite in office areas. Clustering similar nodes and then training the network in groups is a common method used in previous studies [18]. However, it is still a challenge to accurately cluster nodes with similar properties. And separating the network into multiple groups for training makes it difficult to share information between groups and consumes more computational resources and storage space. Dynamic influences and demand data are usually isomorphic, both characterized by changes over time. These dynamics include charging price, charging environment, and other variables that affect charging demand in real time. For example, price influences the charging choices of price-sensitive users, and temperature largely affects the power consumption and charging efficiency of electric vehicles [19]. Previous research mainly used convolution to fuse dynamic external data. Limited by the ambiguity of convolution, these external data bring a certain but weak improvement in prediction accuracy [20, 21, 22]. A more precise method for determining the significance of dynamic influences is needed.
In summary, to build an accurate citywide EV charging demand prediction model that can take into account heterogeneous multi-source influences, several challenges are emerging, which include:
  1. [1)]
  2. 1.
    Static heterogeneous data fusion: Since areas with similar attributes may have similar data variations, mining regional attribute information becomes a promising topic for improving prediction accuracy. A suitable method for clustering areas and effective networks to extract the same patterns of areas with similar land use remain to be explored.
  3. 2.
    Dynamic isomorphic data fusion: Time-series influencing data are fused with demand data mainly in a convolutional approach in previous studies, which results in computational ambiguities and uncertainties in the weights. Different variables have different contribution to prediction and a network that effectively selects and considers variables should be employed to enhance information identification.
  4. 3.
    Structure optimization of prediction network: It remains to be explored how classical neural networks can be improved to more suitably adapt to citywide EV charging demand prediction. For example, possibly not every timestamp contributes sufficiently to the predicted results, and we need to explore sparse and rational allocation of attention in temporal attention networks.
@@
To address the above challenges, we propose an attention-based heterogeneous multivariate data fusion approach for citywide EV charging demand prediction, named AHMDF. It enables 1) classifying areas based on an adaptive clustering method and simultaneously learning similar patterns from areas with similar properties through attentive hypergraph; 2) information propagation between neighboring areas through graph attention; and 3) fusing time-series external variables and efficiently capturing temporal features through multivariate gated Transformer. We demonstrate the state-of-the-art performance of the proposed AHMDF method through a dataset collected from Shenzhen, a first-tier city in China. We also demonstrate the contribution of multiple spatio-temporal influences to accurate EV charging demand prediction through the case study.
@@
  • We propose a novel deep learning framework for citywide electric vehicle charging prediction that can comprehensively capture the spatio-temporal dependencies and effectively consider charging-related dynamic and static influences. We incorporate an area attributes fusion module, an adjacency graph fusion module, and a temporal feature fusion module to achieve accurate prediction.
  • We propose a geo-based clustered hypergraph network for considering urban properties around charging service facilities. We propose an area clustering scheme that combines Term Frequency-Inverse Document Frequency (TF-IDF) and K-means, which can adaptively determine the importance of POI categorization and then cluster urban areas by functions.
  • We propose multivariate gated Transformer to enable the integration of a variable selection network and an improved Transformer encoder. It computes the attention scores of the variables through gated residual networks (GRNs) and efficiently combines external dynamic data. Furthermore, it utilizes GRNs and Gumbel-Softmax to improve the Transformer encoder, which determines sparse attention for the timestamps and avoids the fuzzy perception of long-term attention.
The remaining article is structured as follows. Section 2 reviews the relevant literature. Section 3 presents some preliminary concepts and the problem formulation. Section 4 describes our proposed method, AHMDF, in detail. The method is validated in Section 5 with a representative open access dataset. Section 6 draws conclusions and future directions.
@@ 2 Literature review
Neural networks and deep learning can not only accurately model temporal relationships, but also learn multivariate and multi-node features, which become a research hotspot for citywide electric vehicle charging demand prediction. Recurrent neural networks (including Long Short Term Memory (LSTM)[27, 28] and Gated Recurrent Unit (GRU)[29]) make an important contribution to the accuracy of regression. A study proposes a new hybrid LSTM neural network that contains both historical charging state sequences and time-related features, achieving excellent time-series EV charging demand prediction [9]. Graph Convolutional Neural Network (GCN) is an application of Convolutional Neural Networks (CNN) on graphs, which enables information from neighbors to be learned [11]. The combination of graph neural networks and recurrent neural networks provides the model with powerful spatio-temporal modeling capabilities [15]. A study combining GCN and GRU achieves excellent results on traffic prediction tasks [30]. Moreover, a study improved the spatio-temporal learning capability of the model by replacing the linear operation in GRU with graph convolution [31]. Attention mechanism enables the weights in the model to be determined more rationally and accurately. Transformer can not only determine the importance of temporal feature, but also parallelize the hidden state processing to achieve a faster computation speed than recurrent neural networks [12, 32]. Graph Attention Network (GAT) is a combination of GCN and attention mechanism that allow weights to be determined on graphs [13]. An air traffic density prediction study combines GAT and LSTM to make accurate and robust predictions that outperform baseline models [33]. However, recent research has shown that not every feature contributes sufficiently to the output. Classical attention mechanisms may allocate attention to features that do not contribute enough, thus limiting prediction accuracy. Gated attention network for sequence data is thus proposed [34].
@@
Apart from EV charging demand data, more and more studies are beginning to focus on other influencing factors for assisting in accurate prediction [35, 36, 37]. One of the major influencing factors is the price, which affects the users’ choices. Incorporating price information into the model can dynamically capture user behavior and reduce prediction lag. A deep learning charging demand prediction study utilizes attention mechanisms and bi-directional recurrent neural networks to take into account the effect of price and achieve accurate predictions [38]. Our previous studies use mixed pseudo-sample meta-learning and physics-informed neural networks to model the relationship between charging demand and price, respectively, but these studies still lack consideration of other influencing factors [20, 21]. Moreover, land utilization information (e.g., points of interest) can help determine the amenities surrounding a charging station. A demand prediction study first learns spatio-temporal information by using GCN and GRU, then categorizes EV charging stations based on points of interest, and handles each class of charging stations separately using fully connected networks [18]. However, accurate methods for clustering and fusing intragroup information are still lacking. Day and night show the same cyclical variation as temperature (i.e., high temperatures during the day and low temperatures at night), and EV power consumption is closely related to the environment (e.g. EVs consume more electricity when temperatures are low). In previous studies, environment variables are used to assist in predicting traffic behavior such as parking [35], but there is still a lack of charging demand prediction studies that take environment patterns into account.
Term Frequency-Inverse Document Frequency (TF-IDF) is first used for information retrieval to obtain accurate results in areas such as searching and literature classification [39]. It determines the value of each word and identifies the keywords that are most important to the text. A study presents a novel probabilistic retrieval model, which forms a basis to interpret the TF-IDF term weights as making relevance decisions [40]. This study demonstrates the important role of TF-IDF in searching for critical information. A recent multi-view joint graph learning study applies TF-IDF to identify regions’ intrinsic key attributes and achieves outstanding results [41].
@@
Existing graph neural network frameworks are deployed based on simple graph structures, which limits their application when dealing with complex associations of grouped data. Recently hypergraph-based methods have been proposed to solve this problem [42]. A general high-order multi-modal/multi-type modeling framework called HGNN+ is proposed to simultaneously handle multi-hop adjacencies, attribute groupings, and other hyperedge relations under a single hypergraph-based framework [43]. Hypergraph convolutional neural networks that allow information to propagate over hypergraphs have a wide range of applications in both regression and classification tasks, such as multi-task traffic prediction [44], metro passenger flow prediction [45], POI recommendation [46] and student academic performance classification [47]. Moreover, several studies utilize attention mechanisms to improve hypergraph convolutional networks to select attributes that are important for interpretation [48, 49].
@@
In summary, deep learning enables models with powerful feature extraction capabilities, thus researchers are working on building exquisite and efficient deep learning models for accurate EV charging demand prediction. Recent researches are turning to advanced attention mechanisms to better determine the weights of the network. A growing number of studies consider multi-source factors related to EV charging demand. However, models that can comprehensively and effectively account for both static and dynamic external influences are still lacking. We still need to explore more refined approaches to learn complex spatio-temporal features for accurate citywide EV charging demand prediction.
3 Problem Formulation
First, we divide a city into sub-areas based on travel community for the convenience of urban studies. Electric vehicle charging demand is considered the same as the number of EV charging piles occupied in a given sub-area or travel community. Charging demand in the whole city is defined as $\bm{D}=\{\bm{D}_{:,1},\bm{D}_{:,2},...,\bm{D}_{:,t},...\}\in\mathbb{R}^{N_{A}\times N_{T}}$, where $N_{A}$ is the number of areas, and $N_{T}$ is the number of time steps. Demand at the t-th moment is defined as $\bm{D}_{:,t}=\{D_{1,t},D_{2,t},...,D_{N_{A},t}\}$, where $D_{N,t}$ stands for the charging demand at the $t$-th moment in the $N$-th community. Similar to the demand data, dynamic factors can be defined as $\bm{U}=\{\bm{U}_{:,1},\bm{U}_{:,2},...,\bm{U}_{:,t},...\}\in\mathbb{R}^{N_{A}\times N_{T}}$. Among the dynamic influences, we take price and temperature as examples in our experiments. It is worth noting that our methodology works for other potential dynamics as well, including traffic flow, economic indicators, etc. We slice the data using lookback mechanism with a window size $\tau$. The sliced demand data is denoted as $\bm{D}_{s}=\{\bm{D}_{:,t-\tau+1},\bm{D}_{:,t-\tau+2},...,\bm{D}_{:,t}\}$. The same operation is performed on other multi-source temporal data to get $\bm{U}_{s}$ as inputs of the model. To consider regional associations and further improve prediction accuracy, areas adjacencies (i.e., an undirected graph $\bm{G}$) and area attributes (e.g., number of different types of POIs in the area) are also considered in the model, which are heterogeneous data. Therefore, we can define the citywide electric vehicle charging demand prediction problem that considers urban region and dynamic influences as Eq. (1).
$\left\{\bm{D}_{s},\bm{U}_{s},\bm{G},\rm{POIs}\right\}\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}\bm{\hat{D}}_{t+\tau^{\prime}}$ ((1))
where $\bm{\hat{D}}_{t}$ denotes the predicted electric vehicle charging demand at the $t$-th moment.
@@ 4 Methodology
Fig. 2 illustrates the structure of AHMDF, which consists of three modules, namely a) an area attributes fusion module, in which areas with similar attributes are clustered by using TF-IDF and K-means, and grouped features are learned by using attentive hypergraph neural network; b) an adjacency graph fusion module, which allows for learning adjacency features through graph attention mechanism; and c) a temporal feature fusion module, which enables multi-source temporal variable selection and temporal feature learning through attention mechanisms and improved Transformer encoder.

### v2

@@ Abstract
Electric vehicle charging demand prediction is important for vacant charging pile recommendation and charging infrastructure planning, thus facilitating vehicle electrification and green energy development. The performance of previous spatio-temporal studies is still far from satisfactory nowadays because urban region attributes and multivariate temporal influences are not adequately taken into account. To tackle these issues, we propose a learning approach for citywide electric vehicle charging demand prediction, named CityEVCP. To learn non-pairwise relationships in urban areas, we cluster service areas by the types and numbers of points of interest in the areas and develop attentive hypergraph networks accordingly. Graph attention mechanisms are employed for information propagation between neighboring areas. Additionally, we propose a variable selection network to adaptively learn dynamic auxiliary information and improve the Transformer encoder utilizing gated mechanisms for fluctuating charging time-series data. Experiments on a citywide electric vehicle charging dataset demonstrate the performances of our proposed approach compared with a broad range of competing baselines. Furthermore, we demonstrate the impact of dynamic influences on prediction results in different areas of the city and the effectiveness of our area clustering method.
@@ 1 Introduction
As the world accelerates its transition to low-carbon and environmental friendly mobility, electric vehicles (EVs) will play an important role in the future automotive market [1]. The International Energy Agency points out that at present, the global share of electric vehicles in new car sales is about 20%, and this proportion will rise to about 50% by 2030, while China has already reached such a level in 2024 [2]. However, the lack of available and reasonably located charging facilities is a major constraint to the development of electric vehicles [3]. Providing unlimited charging infrastructure for EV users is impossible in terms of investment costs and energy deployment [4]. The increasing development of smartphones, in-vehicle devices and information distribution systems makes governments and mangers glad to turn to accurate EV charging demand prediction for help. By predicting the services demand in the coming future, administrators can provide users with suggestive information, such as timing and route planning, to avoid customer losses. For governments, accurate charging demand prediction, especially at the citywide level, is a groundwork for a variety of smart strategies such as charging infrastructure planning and smart power supply [5, 6].
Researchers have made many efforts to enable accurate citywide electric vehicle charging demand prediction. Deep learning research provides powerful methods for extracting complex patterns in data variations nowadays. Many studies consider that EV charging demand in future periods is related to historical demand with spatial spillover effects [7]. Therefore, lots of exquisite models introduce Recurrent neural networks to extract temporal features [8] and Graph neural networks to extract spatial features [9], respectively. However, the ambiguity in the weight assignment of convolutional and recurrent operations limits their capabilities. In the latest researches, time-series attention networks (e.g., Transformer) [10] and graph attention networks [11] that are able to better determine the weights on the network have become popular tools for improving the spatio-temporal prediction accuracy. However, previous studies have lacked sufficient consideration of the characteristics of charging demand that are different from other spatio-temporal data [12]. They do not consider that some short periods of fast charging behavior may cause fluctuations in the historical data, which results in not all hidden states can provide enough useful information for prediction. Sparse and rational allocation of attention for accurate charging demand prediction in attention networks remains to be explored.
Furthermore, EV charging behavior is affected by several influencing factors, as shown in Fig. 1. With the application of information collection systems and large-scale data processing centers, various charging-related factors can be effectively centralized and analyzed. Considering adequate factors in urban EV charging demand prediction has become an urgent and promising topic. But there is still a lack of prediction studies that can effectively consider multiple influences at the same time.
@@
Charging-related factors can be categorized into two main groups, namely static ones and dynamic ones. Static influences include neighborhood charging services and land use around the charging stations. These influences are heterogeneous to the time-series charging demands and are usually not directly acceptable as inputs to the network. For example, improvements of the road network and the smart application on mobile devices make it possible for demand in neighboring areas to influence each other. Previous studies employ graph neural networks to fuse information from adjacent nodes and achieve good results [13, 14]. Moreover, some areas have similar land uses and residents may have typical and similar charging habits within those areas [15]. For example, charging demand in commercial areas is relatively high on weekends and after work on weekdays, while it’s the opposite in office areas. Clustering similar nodes and then training the network in separate groups is a common method used in previous studies [16]. However, it is still a challenge to accurately cluster nodes with similar properties. And separating the network into multiple groups for training makes it difficult to share information between groups and consumes more computational resources and storage space.
Dynamic influences and demand data are usually isomorphic, both characterized by changes over time. These dynamics include charging price, charging environment, and other variables that affect charging demand in real time. For example, price influences the charging choices of price-sensitive users, and temperature largely affects the power consumption and charging efficiency of electric vehicles [7]. The relationship between external information and charging demands is often complex and varied in temporal and spatial dimensions. Therefore we need to accurately capture the nonlinearities involved. Previous research mainly used convolution to fuse dynamic external data. Limited by the ambiguity, the convolution methods only bring a certain but weak improvement in prediction accuracy [17, 18]. It is still a challenge to determine the different roles of external auxiliary information in each urban area. A more precise method for determining the citywide significance of dynamic influences is needed.
@@
To address the above challenges, we propose a novel approach for citywide EV charging demand prediction, named CityEVCP, that can effectively consider urban region and dynamic influences simultaneously. It enables 1) classifying areas based on an adaptive clustering method and simultaneously learning similar patterns from areas with similar properties through attentive hypergraph; 2) information propagation between neighboring areas through graph attention; and 3) fusing time-series external variables for city’s each areas through variable selection network and efficiently capturing temporal features through an improved gated Transformer encoder. We demonstrate the state-of-the-art performance of the proposed CityEVCP method through a dataset collected from Shenzhen, a first-tier city in China. We also demonstrate the contribution of multiple influences to accurate EV charging demand prediction through the case study and show that the proposed model can correctly perceive charging-related data patterns.
@@
  • We propose a novel deep learning framework for citywide electric vehicle charging prediction that can effectively consider charging-related dynamic and static influences. We incorporate an area attributes fusion module, an adjacency graph fusion module, and a temporal feature fusion module to achieve accurate prediction.
  • We propose a scheme for considering urban properties around charging service facilities. We view the points of interest (POIs) in an area as a document and propose an adaptive clustering method based on the importance of POIs. Attentive hypergraph mechanism is used to learn common features of charging behavior in grouped urban areas.
  • Considering the different role of dynamics on areas, we utilize gated residual networks (GRNs) to propose a nonlinear variable selection network that adaptively determines the importance of multiple dynamic influences in each urban area. Furthermore, to flexibly adapt to fluctuations in charging demand, we improve the Transformer encoder through GRNs and Gumbel-Softmax, which determines sparse attention for the timestamps and avoids fuzzy perception.
The remaining article is structured as follows. Section 2 reviews the relevant literature. Section 3 describes our proposed method, CityEVCP, in detail. The method is validated in Section 4 with a representative citywide EV charging demand dataset. Section 5 draws conclusions and future directions.
@@ 2 Literature review
Neural networks and deep learning can not only accurately model temporal relationships, but also learn multivariate and multi-node features, which become a research hotspot for citywide electric vehicle charging demand prediction. Recurrent neural networks (including Long Short Term Memory (LSTM)[24, 25] and Gated Recurrent Unit (GRU)[26]) make an important contribution to the accuracy of regression. A study proposes a new hybrid LSTM neural network that contains both historical charging state sequences and time-related features, achieving excellent time-series EV charging demand prediction [8]. Graph Convolutional Neural Network (GCN) is an application of Convolutional Neural Networks (CNN) on graphs, which enables information from neighbors to be learned [9]. The combination of graph neural networks and recurrent neural networks provides the model with powerful spatio-temporal modeling capabilities [13]. A study combining GCN and GRU achieves excellent results on traffic prediction tasks [27]. A study improved the spatio-temporal learning capability of the model by replacing the linear operation in GRU with graph convolution [28]. Attention mechanism enables the weights in the model to be determined rationally and accurately. Transformer can not only determine the importance of temporal feature, but also parallelize the hidden state processing to achieve a faster computation speed than recurrent neural networks [10, 29]. Graph Attention Network (GAT) is a combination of GCN and attention mechanism that allow weights to be determined on graphs [11]. An air traffic density prediction study combines GAT and LSTM to make accurate and robust predictions that outperform baseline models [30]. However, previous research has focused mainly on how to skillfully combine existing tools, while ignoring data characteristics in specialized domains. In the task of EV charging demand prediction, few studies have specifically optimized the network structure in response to the short-term fluctuations in charging demand data.
@@
Apart from EV charging demand data, more and more studies are beginning to focus on other influencing factors for assisting in accurate prediction [31, 32, 33]. One of the major influencing factors is the price, which affects the users’ choices. Incorporating price information into the model can dynamically capture user behavior and reduce prediction lag. A deep learning charging demand prediction study utilizes attention mechanisms and bi-directional recurrent neural networks to take into account the effect of price and achieve accurate predictions [34]. Previous studies use mixed pseudo-sample meta-learning and physics-informed neural networks to model the relationship between charging demand and price, respectively, but these studies still lack consideration of other influencing factors [35, 17]. Day and night show the same cyclical variation as temperature (i.e., high temperatures during the day and low temperatures at night), and EV power consumption is closely related to the environment (e.g. EVs consume more electricity when temperatures are low). In previous studies, environment variables are used to assist in predicting traffic behavior such as parking [31], but there is still few charging demand prediction studies that take environment patterns into account. A study fine-tunes a large language model to enable few-shot predictions of citywide EV charging demand using socio-economic and environmental knowledge pre-trained by the large language model [36].
@@
Land use information (e.g., points of interest) can help determine the amenities surrounding a charging station. A demand prediction study first learns spatio-temporal information by using GCN and GRU, then categorizes EV charging stations based on points of interest, and handles each class of charging stations separately using fully connected networks [16]. However, handling grouped nodes separately is not efficient and hinders some of the information dissemination. Accurate methods for clustering and fusing intragroup information still remains to be explored. A recent multi-view joint graph learning study applies text sequence information retrieval method to identify regions’ intrinsic key attributes and achieves outstanding results [37]. Furthermore, existing graph neural network frameworks are deployed based on simple graph structures for one-by-one paired nodes, which limits their application when dealing with complex associations of grouped data. Recently hypergraph-based methods have been proposed to solve this problem [38]. A general high-order multi-modal/multi-type modeling framework called HGNN+ is proposed to simultaneously handle multi-hop adjacencies, attribute groupings, and other hyperedge relations under a single hypergraph-based framework [39]. Hypergraph convolutional neural networks that allow information to propagate over hypergraphs have a wide range of applications in both regression and classification tasks, such as multi-task traffic prediction [40], metro passenger flow prediction [41], POI recommendation [42] and student academic performance classification [43]. Moreover, several studies utilize attention mechanisms to improve hypergraph convolutional networks to select important attributes [44, 45]. Research on the application of these latest solutions to accurate citywide EV demand prediction remains to be explored.
@@
In summary, deep learning enables models with powerful feature extraction capabilities, thus researchers are working on building exquisite and efficient deep learning models for accurate EV charging demand prediction. Recent researches are turning to advanced attention mechanisms to better determine the weights of the network. A growing number of studies consider multi-source factors related to EV charging demand. However, models that can comprehensively and effectively account for both urban region and dynamic external influences are still lacking. We still need to explore refined approaches to learn complex spatio-temporal features for accurate citywide EV charging demand prediction.
@@ 3 Methodology
We divide a city into sub-areas based on travel community for the convenience of urban studies. Electric vehicle charging demand is considered the same as the number of EV charging piles occupied in a given sub-area or travel community. Fig. 2 illustrates the structure of CityEVCP, which consists of three modules, namely a) an area attributes fusion module, in which areas with similar attributes are clustered according to POI importance, and grouped features are learned by using attentive hypergraph neural network; b) an adjacency graph fusion module, which allows for learning adjacency features through graph attention mechanism; and c) a temporal feature fusion module, which enables multi-source temporal variable selection and temporal feature learning.